OpenStudy (anonymous):

evaluate the function at the given values of x, if possible. f(x)=7x^2-1 over x^2 a. x=3 b. x=-1 Can someone help me solve this

OpenStudy (anonymous):

\[f(3)=\frac{7\times 3^2-1}{3^2}=\frac{7\times 9-1}{9}=\frac{63-1}{9}=\frac{62}{9}\]

OpenStudy (anonymous):

so that is the first one

OpenStudy (anonymous):

yes. for the second one replace x by -1 still need help or you got it?

OpenStudy (anonymous):

no still need help. i dont even know where to begin

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

where you see an x, put in the number

OpenStudy (anonymous):

replace it. so for example of \[f(x)=x^2\] then \[f(3)=3^2\]and if \[f(x)=\frac{x-1}{x}\] then \[f(3)=\frac{3-1}{3}\]

OpenStudy (anonymous):

in your case \[f(x)=\frac{7x^2-1}{x^2}\] so \[f(-1)=\frac{7\times (-1)^2-1}{(-1)^2}=\frac{7\times 1-1}{1}=7-1=6\]

OpenStudy (anonymous):

OMG! This is such jiberish

OpenStudy (anonymous):

Thanks